OpenStudy (anonymous):

Jake's band played songs at three different parties. At the first party they played 4/5 as long as they did at the second party, which was 1/2 as long as the third party. If they played for 6 minutes at the first party, how many minutes did they play at the third party?

OpenStudy (anonymous):

If a,b and c are the lengths of the parties in order, then $$a = \frac{4}{5}b \text{ and } b=\frac{1}{2}c.$$ We can substitute into each other to find out the values since a=6.

OpenStudy (anonymous):

umm idkk lol im stuck on this question??

OpenStudy (anonymous):

Okay so $$6=\frac{4}{5}b \implies b=\frac{6\times 5}{4} = 7.5.$$ Do the same for the next equation :)

OpenStudy (anonymous):

but thats what i came up with but thats not on the answer list

OpenStudy (anonymous):

But it's c that you're interested in.

OpenStudy (anonymous):

correct the answer choices are 2,4 4/5,8 3/5,or 15

OpenStudy (anonymous):

Well b=7.5 and c=2b so c=2 x 7.5 = 15.

OpenStudy (anonymous):

ohhh ohkk i see what i did wrong thank you

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

you guys are so stupid. its A.